Akua is a School of Agriculture graduate and consumes 2 goods: Yoghurt and Bread. Akua also earns a typical student income from the parents, GHS 400 a month. She can either spend it all on Yoghurts and get 50, or she can spend it all on Bread and get 200 loaves.

 

Now assume that Yoghurt and Bread are compliments for Akua. For her to consume all her income of GHS 400, and the prices used in part a. What if Yoghurt is so SOUR that Akua needs 2 loaves of bread for every Yoghurt that she consumes?


1
Expert's answer
2021-05-05T16:41:39-0400

Since the question seems incomplete, i tried understanding the question and thought it should be:

Akua is a School of Agriculture graduate and consumes 2 goods: Yoghurt and Bread. Akua also earns a typical student income from the parents, GHS 400 a month. She can either spend it all on Yoghurts and get 50, or she can spend it all on Bread and get 200 loaves.


(a) Assuming that Yoghurt and Bread are compliments for Akua, construct the equation for Akua’s Budget Line for her to consume all her income of GHS 400.

(b) What if Yoghurt is so SOUR that Akua needs 2 loaves of bread for every Yoghurt that she consumes?


Solution

(a)

Akua’s budget constraint is GHS M : 400

She can either buy 50 yogurt Y

Which means price of each yogurt would be PY: Total budget/ number of yogurts


"=\frac{400}{50}=GHS8"


Or she can buy 200 loaves L of bread with GHS 400


Price of each loaf would be"PL =\frac{ 400}{200} = GHS 2"


The goods are complement her budget constraint would be:


"(PY *\times )Y +( PL \times L) = M"

"8Y + 2L = 400"


(b)

She consumes 2 loaves for every Yoghurt, 

So we can write :"2L = Y"

Putting this value in the budget equation we get:

"8Y + Y = 400"

"9Y = 400"

"Y = 44.444"

"2L = 44.44"

"L = 22.22"


Loaves would be 22.22 and yogurt would be 44.44 to exhaust her budget